Jane went fishing and caught a whopper of a fish! The head of her fish was ten inches long. The length of the fish's tail was equal to the length of the head plus half the length of the body. The body was as long as the head and tail together. What is the total length of her fish?





Answer by rapaljer(4671)

Let x = length of the body, since that is at the end of the sentence.
Now, the tail = length of the head plus half the length of the body.
tail = 10 + 1/2*x

The equation is that the body = head + tail
x= 10 + 10+ 1/2x

Subtract 1/2x from each side:
1/2x= 20

Multiply both sides by 2:
x=40 inches = Body
Tail = 10 + 1/2x
Tail = 10 + 20 = 30 inches

Total fish = Head + Body + Tail
Total fish = 10+40+30 = 80 inches! (BIG FISH!!)

Check:
Body = Head + Tail
40=10+30

It checks!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus